# Melbourne, VIC 3010, Australia

### v2

@@ 1 Introduction
Since its emergence as a theory of quantum gravity, string theory has reached significant milestones, particularly in the precise microscopic counting of black hole entropy Strominger:1996sh ; Maldacena:1997de and the development of the holographic correspondence Maldacena:1997re . The field of holography has expanded dramatically, producing (among much else) numerous formal results that focus on the exact computation of field theory partition functions via supersymmetric localization, Nekrasov:2002qd ; Pestun:2007rz ; Pestun:2016zxk , often linked to branes wrapping cycles on curved manifolds. In the bulk, effective supergravity analysis frequently takes center stage Gauntlett:2003di ; Grana:2005jc , but finite $N$ calculations in field theory are driving us beyond the leading two-derivative approximation. These advancements owe much to the mathematical developments of Duistermaat–Heckman Duistermaat:1982vw and Berline–Vergne–Atiyah–Bott berline1982classes ; Atiyah:1984px in equivariant localization. Parallel to holography, but evolving somewhat independently due to the direct relation to formal mathematics, the pursuit of exact quantum calculations has fueled major progress in topological string theory, Witten:1988xj ; Dijkgraaf:1990dj ; Witten:1992fb ; Bershadsky:1993cx . The development of the topological vertex, Aganagic:2003db ; Aganagic:2003qj , and its refinement on toric Calabi–Yau manifolds, Iqbal:2007ii ; Huang:2010kf ; Krefl:2010fm , have yielded a solution for all string amplitudes. This was eventually applied holographically, resulting in impressive non-perturbative matches with Chern–Simons-matter theories, Aharony:2008ug ; Kapustin:2009kz ; Drukker:2011zy ; Hatsuda:2013oxa . Most recently, lower-dimensional results have sparked interest in the non-perturbative nature of holography and quantum gravity, Saad:2019lba ; Penington:2019npb ; Almheiri:2019psf ; Penington:2019kki ; Almheiri:2019qdq , and have put forward the question of ensemble averages. The present work, along with a companion paper, Cassia:2024 , addresses all the aforementioned topics in a novel manner, aiming to open a new avenue for quantum gravity calculations within string and M-theory.
@@
Figure 1: Conceptual sketch of exact holography from string/M-theory point of view, based on results in Martelli:2023oqk ; Colombo:2023fhu ; Cassia:2024 and here. Blue contours signify a “$\lambda$-ensemble” of equivariant geometric moduli, while yellow correspond to $N$ or $G_{\rm N}$-ensembles in field theory and supergravity, respectively. Green arrows show a change of ensemble, i.e. the relation holds via a forward or inverse transform.
@@
Before quantifying this proposal in Subsection 1.1 and Figure 2, we first give a broader introduction to the main subject of the present paper, equivariant topological string theory. More concretely, as a first step in the present work we focus our attention on the constant maps, i.e. the perturbative part of the topological string partition function encoding the information of classical geometry, and provide explicit results for a set of interesting examples of toric manifolds. These examples also serve as a foundation for holographically dual matches with field theory localization results at finite $N$, which we explore in Cassia:2024 . We have thus divided our work into two parts: the purely mathematical aspects along with the basic holographic setup, presented here; and the applied holographically driven aspects (which for the moment remain less mathematically rigorous), detailed in Cassia:2024 in relation to M2-brane partition functions.
@@
with $t$ the (complexified) Kähler moduli of the manifold $X$, parametrizing the volumes of the two-cycles in $H_{2}(X)$, and $g_{s}$ the string coupling constant. Each of the partition functions $F_{X,\mathfrak{g}}$ in turn enjoys an expansion in terms of the so called Gromov–Witten invariants $N^{X}_{\mathfrak{g},\beta}$ that in a certain sense count the number of holomorphic maps from a genus-$\mathfrak{g}$ Riemann surface into the manifold $X$ [^1] Here, we focus directly on the topological string description through Gromov–Witten theory, rather than the worldsheet perspective. See Givental:1996equ for previous work on equivariant Gromov–Witten theory, which however does not explicitly consider the constant maps terms.
@@
where the power of $Q:=\mathrm{e}^{-t}$ labels the corresponding instanton number (or degree of the map). The leading perturbative contributions to the partition function, corresponding to $\beta=0$ above, are called constant maps terms and enjoy some special geometric interpretation. The genus 00 constant maps term, $N^{X}_{0,0}$, is proportional to the triple intersection numbers, $C_{abc}$, of $X$ and is a homogeneous cubic function of the Kähler moduli, while $N^{X}_{1,0}$ is a linear function proportional to the second Chern numbers of the four cycles of $X$, denoted as $b_{a}$. We come back to the definition of $N^{X}_{\mathfrak{g},0}$ in the next section. More formally, the constant maps terms are determined by the (classical) cohomology ring, while the instanton corrections of the genus 00 Gromov–Witten invariants measure the deformed cohomology ring, or quantum cohomology.
An important observation is that toric Calabi–Yau manifolds, which are necessarily non-compact, allow for a generalization of the topological string partition function via equivariant cohomology.  [^2] Note that toric CY’s are also a natural setting for defining the refined topological string partition function, which we discuss more carefully in Cassia:2024 . Crucially, in the non-compact case, the equivariant upgrade regularizes the volumes of the target manifolds which would otherwise be divergent and in turn it allows for well-defined constant maps terms. The main message here, inspired by the work of Martelli–Zaffaroni and its predecessors, Martelli:2023oqk ; Couzens:2018wnk ; Gauntlett:2018dpc ; Hosseini:2019use ; Hosseini:2019ddy ; Gauntlett:2019roi , is that this has a clear physical meaning. In support of this claim, we can actually consider $X$ to be a Kähler manifold of arbitrary dimension $d$, admitting a torus action $\mathbb{T}^{n}$, with $n$ being the number of homogeneous coordinates. We can thus naturally extend the topological string partition function with a set of equivariant parameters $\epsilon_{i}$, $i=1,\dots,n$. This equivariant generalization is certainly not a novelty in the mathematical physics literature, see Nekrasov:2021ked ; Cassia:2022lfj . It is at the heart of the work in Martelli:2023oqk , who considered the equivariant volume of toric manifolds and its holographic relation to various wrapped brane models, see also Colombo:2023fhu . Importantly, they make use of a set of redundant Kähler parameters, $\lambda^{i}$, that replace the effective Kähler moduli $t^{a}$ (in the language of Ruan:2020qua that we find appropriate to use) which allows us to look at the alternative expansion
@@ 1.1 Main proposal
Having built up some notation, we now state our main conjecture in terms of explicit, albeit still not fully fleshed out, formulae. We consider resolutions, denoted generally by $X$, of the cone over a $(2d-1)$-dimensional Sasakian manifold, $L$, with metric
@@
where $r$ is a coordinate on $\mathbb{R}_{+}$ and $X_{\circ}=C(L)$. The Sasakian condition requires that $X_{\circ}$ is a non-compact Calabi–Yau cone of complex dimension $d$. In the simplest case when $L$ is a round sphere, S2d-1, $X_{\circ}=\mathbb{C}^{d}$ and it does not need to be resolved as it is already smooth. In the generic case $X_{\circ}$ exhibits a conical singularity, and $X$ denotes a smooth resolution of $X_{\circ}$ preserving the CY condition. We further assume it is a toric manifold, such that the Reeb vector field $\xi$ on the space $L$ can be written as
@@
We conjecture that the partition functions of M2-branes with exact charge $N_{\rm M2}$ (related to the brane theory gauge group rank $N$ up to a constant shift), sitting on the tip of the cone $X_{\circ}$ (such that the space $L$ is the near-horizon geometry), is dictated by the full non-perturbative topological string partition function on $X$. Schematically, the $\lambda^{i}$ parameters are dual to a set of fluxes $N_{{\rm M2},i}$, such that the M2-brane partition function on the (possibly squashed) three-sphere, S3, Kapustin:2009kz ; Hama:2011ea ; Imamura:2011wg , is the Laplace transform of the topological string partition function,  [^3] See Section 8 for the description of more general brane systems within string/M-theory.
@@
The expressions are not fully detailed here, as we omit certain subtleties that relate to the mesonic twist (to be defined in Section 2.3) and the topological string refinement (discussed in Cassia:2024 ). However, in essence this proposal represents a quantum generalization of the extremization proposal of Colombo:2023fhu for the $\lambda$ parameters. The extremization corresponds to the saddle point approximation of the integral transform above, including only the leading genus-zero constant maps $N_{0,0}^{X}$ and the leading M2-brane charge, $N$. Consequently, it yields correctly the corresponding large $N$ expression for the partition function $Z_{\rm M2}$ (the right-most part of the holographic map sketched below). We should stress here our 11d point of view, where we regard $g_{s}$ as a formal expansion parameter that has no stringy or geometric origin, but ensures the convergence of the series. In the holographic match it is therefore allowed to be a numerical constant that fixes the normalization.
@@
We defer the explicit details on the integration and the comparison with available M2-brane localization results to Cassia:2024 .
bulkbdry.equivarianttopologicalstrings, $F_{X}(\lambda)$constantmaps terms,$\sum_{\mathfrak{g}}N^{X}_{\mathfrak{g},0}(\lambda)$11d higher der.sugra on $L$, $I_{L}^{\text{HD}}(G_{\mathrm{N}})$11d 2-der.sugra on $L$, $I_{L}(G_{\mathrm{N}})$perturb.trunc.saddle pt.approx.$G_{\mathrm{N}}^{-1}\to\infty$M2-braneson tip of $X$,$Z_{\rm M2}(N_{\rm M2})$$Z^{\text{pert}}_{\rm M2}(N_{\rm M2})$leading freeenergy, ${\cal F}(N)$AdS/CFTperturb.trunc.$N\to\infty$herein Cassia:2024
Figure 2: Schematic diagram of the relations we propose. Blue contour signifies quantities in the $\lambda$-ensemble, while yellow corresponds to the $N$ or $G_{\mathrm{N}}$-ensemble in field theory and supergravity, respectively ($N$ and $G_{\mathrm{N}}$ are related directly via AdS/CFT). Green arrows signify a change of ensemble, and we have suppressed additional indices and (equivariant) parameters.
@@
which we also discuss further and verify at leading order in Cassia:2024  [^4] At this point, it might not be obvious that (9) is on the same footing as (7) from the geometric modeling point of view. We started by defining $X$ as a resolution of $C(L)$, while the fact that $Y$ is a resolution of $C(M)$ follows more subtly from the definition (8). We explore the precise relation between $Y$ and $C(M)$ in Sections 5 and 6 after defining the mesonic twist in Section 2.3.
A specific corollary of the broader conjecture relates to effective supergravity, which serves as an alternative gravitational description of the same physical system. For readers primarily focused on the topological string aspects, the emphasis on the constant maps $N^{X}_{\mathfrak{g},0}(\lambda,\epsilon)$ from the infinite topological string expansion might seem limited. However, our main objective here and in Cassia:2024 is to verify as much as possible the perturbative validity of the general proposal, (7) and (9). Within this scope, the constant maps provide all the essential information. Specifically, we assert that $N^{X}_{0,0}(\lambda,\epsilon)$ fully captures the two derivative supergravity data, while $N^{X}_{1,0}(\lambda,\epsilon)$ (and its refined generalization) encompasses the entire spectrum of higher derivative corrections within supergravity, see Bobev:2021oku ; Hristov:2021qsw ; Hristov:2022lcw . This approach allows us to determine the complete perturbative behavior of any background that effective supergravity can describe.  [^5] Relating the higher-genus constant maps terms $N_{\mathfrak{g}>1,0}$ to supergravity is more subtle and involves a Schiwnger-like calculation, see Dedushenko:2014nya . Note that due to the difference between full integration and saddle point approximation, the higher derivative supergravity terms do not coincide holographically with the perturbative part of the M2-brane partition function, but only to a subset of it. A notable example are the logarithmic corrections, $\log N$, which do follow from the integration of the constant maps but do not correspond to higher derivative corrections in supergravity, see Bhattacharyya:2012ye ; Liu:2017vbl ; Hristov:2021zai ; Bobev:2023dwx .
@@
$F_{X}(\lambda,\epsilon;g_{s})=\log\left(\sum_{N_{\rm M2}}\,\mathrm{e}^{\lambda\,N_{\rm M2}}\,Z^{\rm M2}_{L}(N_{\rm M2},\epsilon)\right)\ .$ (10)
@@
The rest of this paper is organized as follows. In Section 2, we give a brief overview of the topological string partition function and focus on defining the equivariant upgrade of the constant maps terms. In Sections 3, 4, and 5, we provide a comprehensive set of toric examples to build a more hands-on intuition on the subject. While many of these examples have appeared in various forms in other references, our aim is to consolidate them into a uniform notation and focus on some key points. In Section 6, we elaborate further on the relation between the toric manifolds and flux compactifications in string/M-theory. We consider the specific case of vanishing flux backgrounds in Section 7, which provide a special example that completes the results in Sections 35 and relates to the OSV conjecture. We conclude with some general remarks and open directions in Section 8. We leave some additional toric fibration examples for Appendix A as they have no direct flux compactification relevance.

### v4

@@ 1 Introduction
Since its emergence as a theory of quantum gravity, string theory has reached significant milestones, particularly in the precise microscopic counting of black hole entropy Strominger:1996sh ; Maldacena:1997de and the development of the holographic correspondence Maldacena:1997re . The field of holography has expanded dramatically, producing (among much else) numerous formal results that focus on the exact computation of field theory partition functions via supersymmetric localization, Nekrasov:2002qd ; Pestun:2007rz ; Pestun:2016zxk , often linked to branes wrapping cycles on curved manifolds. In the bulk, effective supergravity analysis frequently takes center stage Gauntlett:2003di ; Grana:2005jc , but finite $N$ calculations in field theory are driving us beyond the leading two-derivative approximation. These advancements owe much to the mathematical developments of Duistermaat–Heckman Duistermaat:1982vw and Atiyah–Bott–Berline–Vergne (ABBV) berline1982classes ; Atiyah:1984px in equivariant localization. Parallel to holography, but evolving somewhat independently due to the direct relation to formal mathematics, the pursuit of exact quantum calculations has fueled major progress in topological string theory, Witten:1988xj ; Dijkgraaf:1990dj ; Witten:1992fb ; Bershadsky:1993cx . The development of the topological vertex, Aganagic:2003db ; Aganagic:2003qj , and its refinement on toric Calabi–Yau manifolds, Iqbal:2007ii ; Huang:2010kf ; Krefl:2010fm , have yielded a solution for all string amplitudes. This was eventually applied holographically, resulting in impressive non-perturbative matches with Chern–Simons-matter theories, Aharony:2008ug ; Kapustin:2009kz ; Drukker:2011zy ; Hatsuda:2013oxa . Most recently, lower-dimensional results have sparked interest in the non-perturbative nature of holography and quantum gravity, Saad:2019lba ; Penington:2019npb ; Almheiri:2019psf ; Penington:2019kki ; Almheiri:2019qdq , and have put forward the question of ensemble averages. The present work, along with a companion paper, Cassia:2025jkr , addresses all the aforementioned topics in a novel manner, aiming to open a new avenue for quantum gravity calculations within string and M-theory.
@@
Figure 1: Conceptual sketch of exact holography from string/M-theory point of view, based on results in Martelli:2023oqk ; Colombo:2023fhu ; Cassia:2025jkr and here. Blue contours signify a $\lambda$-ensemble” of equivariant geometric moduli, while yellow correspond to NN or $G_{\rm N}$-ensembles in field theory and supergravity, respectively. Green arrows show a change of ensemble, i.e. the relation holds via a forward or inverse transform.
@@
Before quantifying this proposal in Subsection 1.1 and Figure 2, we first give a broader introduction to the main subject of the present paper, equivariant topological string theory. More concretely, as a first step in the present work we focus our attention on the constant maps, i.e. the perturbative part of the topological string partition function encoding the information of classical geometry, and provide explicit results for a set of interesting examples of toric manifolds. These examples also serve as a foundation for holographically dual matches with field theory localization results at finite $N$, which we explore in Cassia:2025jkr . We have thus divided our work into two parts: the purely mathematical aspects along with the basic holographic setup, presented here; and the applied holographically driven aspects (which for the moment remain less mathematically rigorous), detailed in Cassia:2025jkr in relation to M2-brane partition functions.
@@
with $t$ the (complexified) Kähler moduli of the manifold $X$, parametrizing the volumes of the two-cycles in $H_{2}(X)$, and $g_{s}$ the string coupling constant. Each of the partition functions $F_{X,\mathfrak{g}}$ in turn enjoys an expansion in terms of the so called Gromov–Witten invariants $N^{X}_{\mathfrak{g},\beta}$ that in a certain sense count the number of holomorphic maps from a genus-$\mathfrak{g}$ Riemann surface into the manifold $X$,  [^1] Here, we focus directly on the topological string description through Gromov–Witten theory, rather than the worldsheet perspective. See Givental:1996equ for previous work on equivariant Gromov–Witten theory, which however does not explicitly consider the constant maps terms.
@@
where the power of $Q:=\mathrm{e}^{-t}$ labels the corresponding instanton number (or degree of the map).
The leading perturbative contributions to the partition function, corresponding to $\beta=0$ above, are called constant maps terms and enjoy some special geometric interpretation. These terms correspond to trivial holomorphic maps in Gromov–Witten theory, and we argue they play a key role in linking topological strings to supersymmetric observables. The genus 0 constant maps term, $N^{X}_{0,0}$, is proportional to the triple intersection numbers, $C_{abc}$, of $X$ and is a homogeneous cubic function of the Kähler moduli, while $N^{X}_{1,0}$ is a linear function proportional to the second Chern numbers of the four cycles of $X$, denoted as $b_{a}$. We come back to the definition of $N^{X}_{\mathfrak{g},0}$ in the next section. More formally, the constant maps terms are determined by the (classical) cohomology ring, while the instanton corrections of the genus 0 Gromov–Witten invariants measure the deformed cohomology ring, or quantum cohomology.
An important observation is that toric Calabi–Yau manifolds, which are necessarily non-compact, allow for a generalization of the topological string partition function via equivariant cohomology.  [^2] Note that toric CY’s are also a natural setting for defining the refined topological string partition function, which we discuss more carefully in Cassia:2025jkr . Crucially, in the non-compact case, the equivariant upgrade regularizes the volumes of the target manifolds which would otherwise be divergent and in turn it allows for well-defined constant maps terms. The main message here, inspired by the work of Martelli–Zaffaroni and its predecessors, Martelli:2023oqk ; Couzens:2018wnk ; Gauntlett:2018dpc ; Hosseini:2019use ; Hosseini:2019ddy ; Gauntlett:2019roi , is that this has a clear physical meaning. In support of this claim, we can actually consider $X$ to be a Kähler manifold of arbitrary dimension $d$, admitting a torus action $\mathbb{T}^{n}$, with $n$ being the number of homogeneous coordinates. We can thus naturally extend the topological string partition function with a set of equivariant parameters $\epsilon_{i}$, $i=1,\dots,n$. They correspond to weights of the torus action and regularize infinities in non-compact CY volumes, see Nekrasov:2021ked ; Cassia:2022lfj . This is at the heart of the work in Martelli:2023oqk , who considered the equivariant volume of toric manifolds and its holographic relation to various wrapped brane models, see also Colombo:2023fhu . Importantly, they make use of a set of redundant Kähler parameters, $\lambda^{i}$, that replace the effective Kähler moduli $t^{a}$ (in the language of Ruan:2020qua that we find appropriate to use) which allows us to look at the alternative expansion
@@ 1.1 Main proposal
Having built up some notation, we now state our main conjecture in terms of explicit, albeit still not fully fleshed out, formulae. We consider resolutions, denoted generally by $X$, of the cone over a (2d1)(2d-1)-dimensional toric Sasakian manifold, $L$, with metric
@@
where $r$ is a coordinate on $\mathbb{R}_{+}$ and $X_{\circ}=C(L)$. The toric Sasakian condition requires that $X_{\circ}$ is a non-compact toric local Calabi–Yau cone of complex dimension $d$.  [^3] More precisely, the local Calabi–Yau condition on a non-compact Kähler manifold $X_{\circ}$ requires that the canonical bundle be holomorphically trivial, equivalently that the first Chern class of the tangent bundle vanish in integral cohomology. For conical complex geometries $X_{\circ}=C(L)$, this is equivalent to the Gorenstein condition, namely that the canonical divisor of $X_{\circ}$ is Cartier. The existence of a Sasaki–Einstein metric on $L$ constitutes an additional analytic problem. As discussed in Gauntlett:2006vf , various obstructions to such metrics can occur in general, but for toric Gorenstein cones it is expected that no such obstructions arise, and a Sasaki–Einstein metric can always be realized by an appropriate choice of Reeb vector Martelli:2006yb . In the simplest case when $L$ is a round sphere, S2d-1, $X_{\circ}=\mathbb{C}^{d}$ and it does not need to be resolved as it is already smooth. In the generic case $X_{\circ}$ exhibits a conical singularity, and $X$ denotes a smooth resolution of $X_{\circ}$ preserving the CY condition. The toric assumption means that the Reeb vector field $\xi$ on the space $L$ can be written as
@@
We conjecture that the partition functions of M2-branes with exact charge $N_{\rm M2}$ (related to the brane theory gauge group rank $N$ up to a constant shift, see App. A), sitting on the tip of the cone $X_{\circ}$ (such that the space $L$ is the near-horizon geometry), is dictated by the full non-perturbative topological string partition function on $X$. Schematically, the $\lambda^{i}$ parameters are dual to a set of fluxes $N_{{\rm M2},i}$, such that the M2-brane partition function on the (possibly squashed) three-sphere, S3, Kapustin:2009kz ; Hama:2011ea ; Imamura:2011wg , is the Laplace transform of the topological string partition function,  [^4] See Section 8 for the description of more general brane systems within string/M-theory.
@@
The expressions on the right hand side are not fully detailed here, as we omit certain subtleties that relate to the mesonic twist (to be defined in Section 2.3) and the topological string refinement. A detailed companion work, Cassia:2025jkr , examines both sides of (6) independently and demonstrates perturbative agreement across a range of non-trivial examples, thereby substantiating our proposal. However, in essence this proposal represents a quantum generalization of the extremization proposal of Colombo:2023fhu for the $\lambda$ parameters. The extremization corresponds to the saddle point approximation of the integral transform above, including only the leading genus-zero constant maps $N_{0,0}^{X}$ and the leading M2-brane charge, $N$ (see App. A for the difference between $N_{\rm M2}$ and $N$). Consequently, it yields correctly the corresponding large $N$ expression for the partition function $Z_{\rm M2}$ (the right-most part of the holographic map sketched below). We should stress here our 11d point of view, where we regard $g_{s}$ as a formal expansion parameter that has no stringy or geometric origin, but ensures the convergence of the series. In the holographic match it is therefore allowed to be a numerical constant that fixes the normalization.
@@
We defer the explicit details on the integration and the comparison with available M2-brane localization results to Cassia:2025jkr .
[…]
Figure 2: Schematic diagram of the relations we propose. Blue contour signifies quantities in the $\lambda$-ensemble, while yellow corresponds to the NN or $G_{\mathrm{N}}$-ensemble in field theory and supergravity, respectively (NN and $G_{\mathrm{N}}$ are related directly via AdS/CFT). Green arrows signify a change of ensemble, and we have suppressed additional indices and (equivariant) parameters.
@@
which we also discuss further and verify at leading order in Cassia:2025jkr .  [^5] At this point, it might not be obvious that (9) is on the same footing as (7) from the geometric modeling point of view. We started by defining $X$ as a resolution of $C(L)$, while the fact that $Y$ is a resolution of $C(M)$ follows more subtly from the definition (8). We explore the precise relation between $Y$ and $C(M)$ in Sections 5 and 6 after defining the mesonic twist in Section 2.3.
A specific corollary of the broader conjecture relates to effective supergravity, which serves as an alternative gravitational description of the same physical system. In fact, one can either consider M2-brane backgrounds within 11d supergravity, as done here in Sec. 6 and App. A, or 4d gauged supergravity arising as a consistent truncation on the manifold $L$, as done in our companion paper, Cassia:2025jkr .  [^6] After the first version of this work, reference Gautason:2025plx proposed a further refinement on the supergravity side of the diagram below. The authors of Gautason:2025plx argue that in fact 11d supergravity observables naturally come in the $\lambda$ (blue) ensemble, while 4d supergravity observables are in the $G_{\mathrm{N}}$ (yellow) ensemble. For readers primarily focused on the topological string aspects, the emphasis on the constant maps $N^{X}_{\mathfrak{g},0}(\lambda,\epsilon)$ from the infinite topological string expansion might seem limited. However, our main objective here and in Cassia:2025jkr is to verify as much as possible the perturbative validity of the general proposal, (7) and (9). Within this scope, the constant maps provide all the essential information. Specifically, $N^{X}_{0,0}(\lambda,\epsilon)$ fully captures the two derivative supergravity data (either in 11d or in 4d), as discussed at the level of Sasakian geometry in Martelli:2005tp ; Butti:2005vn ; Butti:2005ps ; Martelli:2006yb ; Amariti:2011uw ; Couzens:2018wnk ; Gauntlett:2018dpc ; Hosseini:2019use ; Hosseini:2019ddy ; Gauntlett:2019roi ; Gauntlett:2019pqg ; Boido:2022mbe and reviewed in Sec. 6; while we assert that $N^{X}_{1,0}(\lambda,\epsilon)$ (and its refined generalization) encompasses the entire spectrum of higher-derivative (HD) corrections within effective 4d supergravity, see Bobev:2020egg ; Bobev:2021oku ; Hristov:2021qsw ; Hristov:2022lcw . Alternatively, from 11d perspective, our conjecture is that all HD corrections come into the (refined) genus-one constant maps, except for the topological CS-term explicitly considered in App. A responsible for the shift in the exact M2-brane charge, $N_{\rm M2}$.  [^7] The difference between the 11d and 4d higher-derivative corrections is simply due to the fact that $N_{\rm M2}$, subject to 11d HD corrections, enters in the value of the effective 4d Newton constant, $G_{\rm 4d}$, see Sec. 6. Thus only the effective 4d supergravity fits clearly in the yellow ensemble, see the previous footnote. This approach allows us to determine the complete perturbative behavior of any background that effective supergravity can describe.  [^8] Relating the higher-genus constant maps terms $N_{\mathfrak{g}>1,0}$ to supergravity is more subtle and involves a Schiwnger-like calculation, see Dedushenko:2014nya . Note that due to the difference between full integration and saddle point approximation, the HD supergravity terms do not coincide holographically with the perturbative part of the M2-brane partition function, but only to a subset of it. A notable example are the logarithmic corrections, $\log N$, which do follow from the integration of the constant maps but do not correspond to HD corrections in supergravity, see Bhattacharyya:2012ye ; Liu:2017vbl ; Hristov:2021zai ; Bobev:2023dwx .
@@
$F_{X}(\lambda,\epsilon;g_{s})=\log\left(\sum_{N_{\rm M2}}\,\mathrm{e}^{\lambda\,N_{\rm M2}}\,Z^{\rm M2}_{L}(N_{\rm M2},\epsilon)\right)\ ,$ (10)
which formally represents the inverse Laplace transform of (6).
@@
The rest of this paper is organized as follows. In Section 2, we give a brief overview of the topological string partition function and focus on defining the equivariant upgrade of the constant maps terms. In Sections 3, 4, and 5, we provide a comprehensive set of toric examples to build a more hands-on intuition on the subject. While many of these examples have appeared in various forms in other references, our aim is to consolidate them into a uniform notation and focus on some key points. In Section 6, we elaborate further on the relation between the toric manifolds and flux compactifications in string/M-theory. We consider the specific case of vanishing flux backgrounds in Section 7, which provide a special example that completes the results in Sections 35 and relates to the OSV conjecture. We conclude with some general remarks and open directions in Section 8. In Appendix A (to be read after Section 6) we calculate the exact M2-brane charge, denoted above by $N_{\rm M2}$. Finally, some additional toric fibration examples are relegated to Appendix B as they have no direct flux compactification relevance.